$SPAD/src/input sae.input

Waldek Hebisch December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)spool sae.output
)set message test on
)set message auto off
)clear all
--S 1 of 6
pol1:=x^2+1
--R
--R         2
--R   (1)  x  + 1
--R                                                    Type: Polynomial(Integer)
--E 1
--S 2 of 6
pol2:=z^3-2
--R
--R         3
--R   (2)  z  - 2
--R                                                    Type: Polynomial(Integer)
--E 2
--S 3 of 6
primrec:=primitiveElement([pol1,pol2],[x,z])$PrimitiveElement(FRAC(INT))
--I
--I   (3)
--I   [coef= [- 3,- 1],
--I
--I     poly =
--I            2   5     1   4    20  3    13   2   2431      91
--I       [- ---- ?  - ---- ?  - --- ?  - ---- ?  - ---- ? + ---,
--I          1293      7758      431      1293      3879     862
--I         2   5     1   4    60  3    13  2   1138     273
--I        --- ?  + ---- ?  + --- ?  + --- ?  + ---- ? - ---]
--I        431      2586      431      431      1293     862
--I     ,
--I           6      4     3       2
--I    prim= ?  + 27?  + 4?  + 243?  - 108? + 733]
--IType: Record(coef: List(Integer),poly: ...
--E 3
--S 4 of 6
Ae:=SAE(FRAC(INT),SparseUnivariatePolynomial(FRAC(INT)),primrec.prim)
--I
--I   (4)
--I  SimpleAlgebraicExtension(Fraction Integer,SparseUnivariatePolynomial Fraction
--I   Integer,?**6+27*?**4+4*?**3+243*?*?+(-108*?)+733)

--R                                                                 Type: Domain
--E 4
--S 5 of 6
(primrec.poly.1::Ae)^2
--R
--R   (5)  - 1
--IType: SimpleAlgebraicExtension(Fraction(Integer),...
--E 5
--S 6 of 6
(primrec.poly.2::Ae)^3
--R
--R   (6)  2
--IType: SimpleAlgebraicExtension(Fraction(Integer),...
--E 6
)spool
)lisp (bye)

———-

References^{[1] nothing}

5